# Contrastive Prompt Clustering for Weakly Supervised Semantic Segmentation

### v1

@@ 1 Introduction
Weakly supervised semantic segmentation (WSSS) utilizes image-level labels to achieve dense pixel-level predictions Wu et al. [2025b, 2024a]. In particular, WSSS refers to methods that rely solely on image-level annotations instead of detailed pixel-level labels. This approach significantly reduces the annotation cost and effort, while still producing high-resolution segmentation maps in which each pixel is assigned a semantic label.
@@
Recently, many WSSS methods Kweon & Yoon [2024], Yin et al. [2023], Chen et al. [2024], Li et al. [2023] adopt the Vision Transformer (ViT) architecture Dosovitskiy et al. [2020], Xue et al. [2025], Wu et al. [2024b], due to its capability to capture global contextual dependencies through self-attention mechanisms. These mechanisms allow ViT to model long-range interactions across an image, thereby improving semantic coherence. However, the self-attention in ViT also tends to act as a low-pass filter, diminishing the variance of local features and leading to excessive smoothing Yin et al. [2024]. This effect often suppresses high-frequency details that are critical for preserving fine-grained boundaries and intra-object distinctions, which ultimately weakens the discriminability of patch-level features and harms segmentation quality. In addition, existing WSSS approaches Li et al. [2024], Yin et al. [2024], Guo et al. [2024] typically emphasize maximizing inter-class separation to reduce class confusion. While this strategy can enhance overall discriminative power, it often overlooks the valuable shared semantic information among visually or semantically similar categories. We argue that mining and leveraging such shared information is essential for capturing the latent structure of the semantic space. To this end, we propose to group related classes into semantic clusters, and inject their shared features into patch representations to mitigate the over-smoothing effect and reduce ambiguity, as illustrated in Fig. 1.
To effectively identify and exploit these inter-category relationships, we introduce an approach called Contrastive Prompt Clustering (CPC). Our method leverages the capabilities of Large Language Models (LLMs) to automatically generate prompts that encode semantic affinities between categories, enabling the construction of prompt-guided clusters. These clusters serve as a form of high-level relational prior, guiding the network to recognize shared information across similar classes.
Furthermore, to strengthen class-level discriminability, we introduce a class-aware contrastive learning module that operates at the patch level. This module enforces intra-class compactness and inter-class separability by pulling embeddings from the same class closer together and pushing embeddings from different classes farther apart. Unlike prompt clustering, which models coarse-grained semantic similarities, the contrastive objective imposes fine-grained visual constraints, even for classes that belong to the same semantic cluster. This hierarchical design enables the model to benefit from both shared semantic priors and class-specific visual distinctions.
@@ 2.1 WSSS with Image-level Labels
Weakly Supervised Semantic Segmentation (WSSS) with image-level labels generally begins by generating pseudo labels using CAM Wu et al. [2024c, 2025a], Zhang et al. [2025], Wu et al. [2025c]. However, since CAMs tend to highlight only the most discriminative regions of an object, numerous methods have been developed to overcome this limitation. For instance, techniques such as erasing Wei et al. [2017], Jiang et al. [2022], online attention accumulation Jiang et al. [2019], and cross-image semantic mining Sun et al. [2020] have been proposed to broaden the activated regions. Additional works Lee et al. [2021], Yao et al. [2021] employ supplementary saliency maps to both suppress background interference and identify non-salient object parts. Other approaches Chen et al. [2022], Du et al. [2022], Cao & Zhang [2023], Yaganapu & Kang [2024], Xu et al. [2019], Luo et al. [2022] leverage contrasts between pixel and prototype representations to encourage more complete object activation. Furthermore, the method in Chang et al. [2020] enhances the segmentation process by integrating extra category information or by refining the learned representations through additional information mining on the training data. Recent advances have also seen the adoption of ViT for WSSS. Models such as MCTformer Xu et al. [2022] and AFA Ru et al. [2022] use ViT’s attention mechanisms to generate localization maps and produce pseudo labels via PSA Ahn & Kwak [2018], and AFA through a combination of multi-head self-attention and an affinity module. Meanwhile, ViT-PCM Rossetti et al. [2022] explores a CAM-independent approach by utilizing patch embeddings and max pooling to estimate pixel-level label probabilities, albeit with the risk of patch misclassification. In contrast to these strategies, our proposed CPC exploits additional cluster information to further enhance the segmentation performance of WSSS by improving its ability to distinguish between different classes.
@@ 2.2 Vision Transformers for WSSS
ViTDosovitskiy et al. [2020] has achieved impressive results across a range of vision tasksXie et al. [2024b], Xu et al. [2019], Luo et al. [2022], Liu et al. [2025, 2024], Xie et al. [2024a], Dosovitskiy et al. [2020]. This success has naturally led to the integration of ViT into WSSS, where ViT-based approaches are emerging as viable alternatives to traditional CNN-based methods for generating CAMXu et al. [2022], Ru et al. [2022]. For example, MCTformer Xu et al. [2022] utilizes the inherent attention mechanisms of ViT to derive localization maps and employs PSA Ahn & Kwak [2018] for pseudo mask generation. Similarly, AFA Ru et al. [2022] capitalizes on ViT’s multi-head self-attention to capture global context and uses an affinity module to propagate pseudo masks. Despite these advances, the challenge of over-smoothing inherent in ViT remains a significant hurdle. ViT-PCM Rossetti et al. [2022] attempts to address this by abandoning CAM in favor of a method based on patch embeddings combined with max pooling to infer pixel-level label probabilities, although this may introduce errors when patches are misclassified. Our approach sets itself apart by incorporating additional cluster information to augment the WSSS framework’s capacity to recognize and differentiate among various classes, thereby addressing some of the limitations present in existing ViT-based methods.
@@ 2.3 Prompt-based Language Models
Prompt-based learning improves the performance of pre-trained language models (PLMs) by incorporating task-specific instructions into the input. Early methods relied on manually designed prompts to adapt to various generation tasks Brown et al. [2020], Raffel et al. [2020], Zou et al. [2021], Chen et al. , Fang et al. [2025], Jin et al. [2025]. However, these handcrafted prompts lack the flexibility to be easily applied to new tasks. As a result, recent advancements have focused on automating prompt generation Shin et al. [2020]. Additionally, some approaches Liu et al. [2023], Li & Liang [2021] have explored optimizing continuous prompts to address the challenges posed by optimizing discrete prompts, offering improved flexibility across different tasks. More recently, PLMs have been applied to computer vision tasks, with one method Zhang et al. [2023] utilizing prompts to enhance few-shot learning. In contrast, our approach leverages PLMs to generate prompts that enhance the diversity of image description text. To the best of our knowledge, this is the first time PLMs have been used to generate cluster information to improve the WSSS task.
@@ 3.1 Overall Framework
As illustrated in Fig. 2, we provide a comprehensive overview of our proposed CPC framework. The network is designed to predict patch-level semantic labels by injecting both visual features and semantic cluster priors into the representation learning process. Specifically, the original ViT patch tokens are concatenated with a set of cluster tokens derived from LLM-guided prompt clustering, resulting in enriched patch representations. These tokens are further refined using a HV-BiLSTM module Rossetti et al. [2022] to better model inter-patch dependencies.
A lightweight MLP-based classifier then predicts the patch-to-category associations, generating coarse pseudo-labels. These initial predictions are post-processed with a Conditional Random Field (CRF) Krähenbühl & Koltun [2011] to enforce spatial consistency and semantic alignment, resulting in high-quality pseudo-labels.
@@
Finally, the refined pseudo-labels are used to train a DeepLabv2 Chen et al. [2018] segmentation network in a fully supervised manner, which produces the final dense semantic predictions. This two-stage training pipeline enables our framework to effectively leverage both language-guided clustering and feature-level supervision, achieving superior performance in weakly supervised semantic segmentation.
@@ 3.2 Self-Refine Prompt for Cluster
The motivation behind clustering categories arises from our observation that, in previous works, models are typically designed to maximize the distance between different classes while minimizing the distance within the same class. This approach often leads to the assumption that the distances between all classes are equidistant, ignoring their intrinsic semantic similarities. For instance, a model might perceive the distance from “cat” to “dog” as equal to the distance from “cat” to “car,” despite the fact that “cat” and “dog” share significantly more similarities as animals, whereas ”cat” and ”car” are entirely distinct objects. To address this limitation and better capture the nuanced relationships between classes, we propose to abstract more granular cluster information at the class level. Specifically, we harness the powerful capabilities of LLMs to assign distinct cluster tags to these classes. However, due to the inherent instability of LLMs Madaan et al. [2024], the generated cluster tags may lack sufficient reliability.
@@ 4 Experiments
In this section, we present our experimental settings and evaluation protocols. We compare our proposed method, Contrastive Prompt Clustering (CPC), with state-of-the-art weakly supervised semantic segmentation (WSSS) approaches on the PASCAL VOC 2012 Everingham et al. [2010] and MS COCO 2014 Lin et al. [2014] benchmarks. Finally, we perform ablation studies to validate the contributions of each component in our framework.
@@ 4.1 Experimental Settings
Datasets and Evaluation Metric. We evaluate the effectiveness of the proposed framework on two widely adopted benchmarks for weakly supervised semantic segmentation: PASCAL VOC 2012 Everingham et al. [2010] and MS COCO 2014 Lin et al. [2014]. The PASCAL VOC 2012 dataset comprises 20 foreground object categories and one background class. Following standard practice, we augment the original training set with additional annotations from the Semantic Boundaries Dataset (SBD) Hariharan et al. [2011], resulting in 10,582 training images annotated with image-level labels and 1,449 images reserved for validation. MS COCO 2014 provides a more challenging evaluation setting, containing 81 object categories with significant intra-class variation and complex scene compositions. In accordance with prior WSSS literature, we employ approximately 82,000 images with image-level supervision for training and 40,000 images from the validation set for performance evaluation.
@@
Implementation Details. In our experiments, we utilize the ViT-B/16 model as the encoder, owing to its demonstrated ability to capture rich visual representations. Each image is resized to $384\times 384$ during training Kolesnikov & Lampert [2016] to ensure consistent input resolution and is subsequently divided into $24\times 24$ non-overlapping patches of size $16\times 16$. This patch-wise representation enables the transformer architecture to model both local and global contextual information through self-attention mechanisms.
@@
For inference, input images are upsampled to $960\times 960$ to generate high-resolution predictions. The patch classifier produces class probability maps, which are subsequently refined using Conditional Random Fields (CRF) Krähenbühl & Koltun [2011] to improve spatial coherence and boundary accuracy. These refined pseudo-labels are employed to supervise a DeepLab V2 Chen et al. [2018] segmentation model, which leverages atrous convolutions to capture multi-scale contextual dependencies, thereby producing the final dense semantic segmentation output.
@@
method Pub. Backbone mIoU AdvCAM Lee et al. [2022] PAMI22 V2-RN101 69.9 SIPE Chen et al. [2022] CVPR22 ResNet50 58.6 AFA Ru et al. [2022] CVPR22 MiT-B1 66.0 ViT-PCM Dosovitskiy et al. [2020] ECCV22 ViT-B/16 71.4 ToCo Ru et al. [2023] CVPR23 ViT-B/16 72.2 USAGE Peng et al. [2023] ICCV23 ResNet38 72.8 FPR Chen et al. [2023] ICCV23 ResNet38 68.5 ToCo Ru et al. [2023] CVPR23 ViT-B/16 70.5 SFC Zhao et al. [2024] AAAI24 ViT-B/16 73.7 APC Wu et al. [2025b] EAAI25 ViT-B/16 74.6 CPC Ours ViT-B/16 75.7
@@ 4.2 Comparison with State-of-the-arts
In addition to the inter-class clustering, our framework integrates a patch-level contrastive loss to further enhance intra-class consistency, which contributes to more coherent and robust pixel-level predictions. Together, these components allow the model to more accurately distinguish between ambiguous regions and reduce misclassification errors caused by overlapping semantics. As shown in Tab. 1, our method achieves superior mIoU performance compared to a range of recent state-of-the-art methods, including both convolution-based and transformer-based backbones. Notably, CPC outperforms strong baselines such as ToCo Ru et al. [2023] and SFC Zhao et al. [2024], demonstrating the effectiveness of our dual-level learning strategy in producing high-quality pseudo labels.
@@
Model Pub. Backbone Val USAGE Peng et al. [2023] ICCV23 ResNet38 71.9 SAS Kim et al. [2023] AAAI23 ViT-B/16 69.5 MCTformer Xu et al. [2022] CVPR22 DeiT-S 61.7 SIPE Chen et al. [2022] CVPR22 ResNet50 58.6 ViT-PCM Dosovitskiy et al. [2020] ECCV22 ViT-B/16 69.3 AFA Ru et al. [2022] CVPR22 MiT-B1 63.8 ToCo Ru et al. [2023] CVPR23 ViT-B/16 70.5 TSCD Xu et al. [2023] AAAI23 MiT-B1 67.3 FPR Chen et al. [2023] ICCV23 ResNet38 70.0 SFC Zhao et al. [2024] AAAI24 ViT-B/16 71.2 IACD Wu et al. [2024a] ICASSP24 ViT-B/16 71.4 APC Wu et al. [2025b] EAAI25 ViT-B/16 72.3 CPC Ours ViT-B/16 73.4
@@
Model Pub. Backbone mIoU (%) MCTformer Xu et al. [2022] CVPR22 Resnet38 42.0 ViT-PCM Dosovitskiy et al. [2020] ECCV22 ViT-B/16 45.0 SIPE Chen et al. [2022] CVPR22 Resnet38 43.6 TSCD Xu et al. [2023] AAAI23 MiT-B1 40.1 SAS Kim et al. [2023] AAAI23 ViT-B/16 44.5 FPR Chen et al. [2023] ICCV23 ResNet38 43.9 ToCo Ru et al. [2023] CVPR23 ViT-B 42.3 SFC Zhao et al. [2024] AAAI24 ViT-B/16 44.6 IACD Wu et al. [2024a] ICASSP24 ViT-B/16 44.3 PGSD Hao et al. [2024] TCSVT24 ViT-B/16 43.9 APC Wu et al. [2025b] EAAI25 ViT-B/16 45.7 CPC Ours ViT-B/16 46.7
Semantic Segmentation Results. To evaluate the effectiveness of our proposed pseudo-label generation strategy, we adopt DeepLab V2 Chen et al. [2018] as the segmentation backbone and train it using the pseudo labels generated by our CPC framework. We report semantic segmentation performance on the PASCAL VOC 2012 validation set and compare it with recent state-of-the-art WSSS approaches, as presented in Tab. 2. Our CPC framework yields superior segmentation performance, achieving an mIoU of 72.2% and outperforming prior transformer-based methods such as ToCo Ru et al. [2023], SFC Zhao et al. [2024], and IACD Wu et al. [2024a]. These results confirm the advantage of combining inter-class category clustering with intra-class contrastive learning for generating high-quality pseudo labels. The enhanced quality of pseudo labels leads to clearer delineation of complex object structures and reduced confusion between visually similar classes.

### v2

@@ 1 Introduction
Weakly supervised semantic segmentation (WSSS) aims to produce pixel-level semantic predictions using only image-level annotations [1, 2, 3, 4]. This paradigm greatly reduces the cost and labor associated with fine-grained labeling, making it an appealing alternative to fully supervised methods. Despite the weak supervision, WSSS methods strive to generate high-resolution segmentation masks where each pixel is assigned an appropriate class label.
@@
Recently, many WSSS methods [5, 6, 7, 8, 9] adopt the Vision Transformer (ViT) architecture [10, 11, 12], due to its capability to capture global contextual dependencies through self-attention mechanisms. These mechanisms allow ViT to model long-range interactions across an image, thereby improving semantic coherence. However, the self-attention in ViT also tends to act as a low-pass filter, diminishing the variance of local features and leading to excessive smoothing [13]. This effect often suppresses high-frequency details that are critical for preserving fine-grained boundaries and intra-object distinctions, which ultimately weakens the discriminability of patch-level features and harms segmentation quality. In addition, existing WSSS approaches [14, 13, 15] typically emphasize maximizing inter-class separation to reduce class confusion. While this strategy can enhance overall discriminative power, it often overlooks the valuable shared semantic information among visually or semantically similar categories. We argue that mining and leveraging such shared information is essential for capturing the latent structure of the semantic space. To this end, we propose to group related classes into semantic clusters, and inject their shared features into patch representations to mitigate the over-smoothing effect and reduce ambiguity, as illustrated in Fig. 1.
To effectively identify and exploit inter-category relationships, we propose Contrastive Prompt Clustering (CPC). Our method leverages Large Language Models (LLMs) to generate prompts that encode semantic affinities between categories, forming prompt-guided clusters that serve as high-level relational priors to guide recognition of shared semantics across similar classes. To further enhance class-level discriminability, we introduce a class-aware patch-level contrastive module that enforces intra-class compactness and inter-class separability, pulling embeddings of the same class closer while pushing apart those of different classes. While prompt clustering captures coarse-grained semantic similarities, the contrastive objective imposes fine-grained visual constraints, even within semantically related clusters. This hierarchical design allows the model to jointly benefit from shared semantic priors and class-specific distinctions.
@@ 2.1 WSSS with Image-level Labels
Weakly Supervised Semantic Segmentation (WSSS) that relies solely on image-level annotations often initiates its training by generating pseudo masks derived from Class Activation Maps (CAM) [16, 17, 18, 19]. Nevertheless, since CAMs typically capture only the most salient object parts, various efforts have emerged to overcome this limitation. To enrich the activated regions, techniques such as erasing [20, 21], attention accumulation during training [22], and semantic consistency mining across images [23] have been proposed. Additionally, some approaches [24, 25] incorporate saliency guidance to suppress irrelevant background and reveal under-highlighted object parts. Other solutions [26, 27, 28, 29, 30, 31] focus on contrastive objectives between pixel embeddings and class prototypes to encourage complete region coverage. Moreover, the work in [32] enhances segmentation by injecting auxiliary category knowledge and refining the learned features through mining label priors in the training set.
Recent methods have also integrated Vision Transformers (ViT) into WSSS pipelines. For example, MCTformer [33] and AFA [34] leverage ViT’s self-attention to generate localization cues, where MCTformer uses PSA [35] and AFA adds affinity-aware refinement modules. In contrast, ViT-PCM [36] discards CAM entirely and estimates pixel-level labels via patch embeddings with max pooling, although it may suffer from patch-level noise. Unlike these paradigms, our CPC framework introduces additional clustering cues to guide the segmentation process, allowing more robust differentiation between semantically related classes and improving overall performance.
@@ 2.2 Vision Transformers for WSSS
Vision Transformer (ViT) [10] has demonstrated strong performance across various vision-related benchmarks [37, 30, 31, 38, 39, 40, 10]. This progress has encouraged its adoption in the field of WSSS, where ViT-based methods are gradually replacing traditional CNN-based CAM generation techniques [33, 34]. For instance, MCTformer [33] utilizes ViT’s self-attention to construct class localization maps and applies PSA [35] to generate pseudo masks. AFA [34] extends this idea by incorporating a multi-head self-attention module alongside an affinity propagation unit to strengthen mask quality through contextual reasoning. Nevertheless, ViT’s tendency towards over-smoothing continues to limit its ability to preserve fine details. In response, ViT-PCM [36] proposes to bypass CAM generation entirely by relying on patch embeddings and max pooling to estimate dense pixel predictions, though it risks introducing noise due to potential patch misclassification. Unlike these prior strategies, our method leverages additional cluster-level information to improve class discrimination, thereby enhancing the WSSS pipeline’s ability to handle visually similar categories and overcoming several limitations inherent to ViT-based segmentation methods.
@@ 2.3 Prompt-based Language Models
Prompt-based strategies have become a powerful mechanism to adapt pre-trained language models (PLMs) to downstream tasks by embedding task-specific cues directly into the input. Initial approaches often relied on manually crafted prompts tailored to particular generation objectives [41, 42, 43, 44, 45, 46]. However, such manually designed prompts tend to suffer from limited generalizability and poor adaptability to unseen tasks. To overcome these issues, subsequent research introduced automatic prompt construction techniques [47], which eliminate the need for manual engineering. Moreover, other methods have focused on continuous prompt tuning [48, 49], which replaces discrete token sequences with trainable embeddings, offering more flexible and effective adaptation across diverse tasks. Recently, the scope of prompt-based PLMs has expanded into the vision domain—for example, [50] demonstrates the benefit of prompts for few-shot learning in visual recognition. Differing from these studies, our method employs PLMs to synthesize diverse textual prompts aimed at enriching image descriptions. To our knowledge, this is the first work that exploits PLM-generated prompts to derive semantic cluster information specifically designed to enhance performance in the WSSS setting.
@@ 3.1 Overall Framework
As illustrated in Fig. 2, we provide a comprehensive overview of our proposed CPC framework. The network is designed to predict patch-level semantic labels by injecting both visual features and semantic cluster priors into the representation learning process. Specifically, the original ViT patch tokens are concatenated with a set of cluster tokens derived from LLM-guided prompt clustering, resulting in enriched patch representations. These tokens are further refined using a HV-BiLSTM module [36] to better model inter-patch dependencies.
A lightweight MLP-based classifier then predicts the patch-to-category associations, generating coarse pseudo-labels. These initial predictions are post-processed with a Conditional Random Field (CRF) [51] to enforce spatial consistency and semantic alignment, resulting in high-quality pseudo-labels.
@@
Finally, the refined pseudo-labels are used to train a DeepLabv2 [52] segmentation network in a fully supervised manner, which produces the final dense semantic predictions. This two-stage training pipeline enables our framework to effectively leverage both language-guided clustering and feature-level supervision, achieving superior performance in weakly supervised semantic segmentation.
@@ 3.2 Self-Refine Prompt for Cluster
The motivation behind clustering categories arises from our observation that, in previous works, models are typically designed to maximize the distance between different classes while minimizing the distance within the same class. This approach often leads to the assumption that the distances between all classes are equidistant, ignoring their intrinsic semantic similarities. For instance, a model might perceive the distance from “cat” to “dog” as equal to the distance from “cat” to “car,” despite the fact that “cat” and “dog” share significantly more similarities as animals, whereas ”cat” and ”car” are entirely distinct objects. To address this limitation and better capture the nuanced relationships between classes, we propose to abstract more granular cluster information at the class level. Specifically, we harness the powerful capabilities of LLMs to assign distinct cluster tags to these classes. However, due to the inherent instability of LLMs [53], the generated cluster tags may lack sufficient reliability.
@@ 4 Experiments
In this section, we present our experimental settings and evaluation protocols. We compare our proposed method, Contrastive Prompt Clustering (CPC), with state-of-the-art weakly supervised semantic segmentation (WSSS) approaches on the PASCAL VOC 2012 [54] and MS COCO 2014 [55] benchmarks. Finally, we perform ablation studies to validate the contributions of each component in our framework.
@@ 4.1 Experimental Settings
Datasets and Evaluation Metric. We evaluate the effectiveness of the proposed framework on two widely adopted benchmarks for weakly supervised semantic segmentation: PASCAL VOC 2012 [54] and MS COCO 2014 [55]. The PASCAL VOC 2012 dataset comprises 20 foreground object categories and one background class. Following standard practice, we augment the original training set with additional annotations from the Semantic Boundaries Dataset (SBD) [56], resulting in 10,582 training images annotated with image-level labels and 1,449 images reserved for validation. MS COCO 2014 provides a more challenging evaluation setting, containing 81 object categories with significant intra-class variation and complex scene compositions. In accordance with prior WSSS literature, we employ approximately 82,000 images with image-level supervision for training and 40,000 images from the validation set for performance evaluation.
@@
Implementation Details. In our experiments, we utilize the ViT-B/16 model as the encoder, owing to its demonstrated ability to capture rich visual representations. Each image is resized to $384\times 384$ during training [57] to ensure consistent input resolution and is subsequently divided into $24\times 24$ non-overlapping patches of size $16\times 16$. This patch-wise representation enables the transformer architecture to model both local and global contextual information through self-attention mechanisms.
@@
For inference, input images are upsampled to $960\times 960$ to generate high-resolution predictions. The patch classifier produces class probability maps, which are subsequently refined using Conditional Random Fields (CRF) [51] to improve spatial coherence and boundary accuracy. These refined pseudo-labels are employed to supervise a DeepLab V2 [52] segmentation model, which leverages atrous convolutions to capture multi-scale contextual dependencies, thereby producing the final dense semantic segmentation output.
@@
method Pub. Backbone mIoU AdvCAM [58] PAMI22 V2-RN101 69.9 SIPE [26] CVPR22 ResNet50 58.6 AFA [34] CVPR22 MiT-B1 66.0 ViT-PCM [10] ECCV22 ViT-B/16 71.4 ToCo [59] CVPR23 ViT-B/16 72.2 USAGE [60] ICCV23 ResNet38 72.8 FPR [61] ICCV23 ResNet38 68.5 ToCo [59] CVPR23 ViT-B/16 70.5 SFC [62] AAAI24 ViT-B/16 73.7 APC [1] EAAI25 ViT-B/16 74.6 CPC Ours ViT-B/16 75.7
@@ 4.2 Comparison with State-of-the-arts
In addition to the inter-class clustering, our framework integrates a patch-level contrastive loss to further enhance intra-class consistency, which contributes to more coherent and robust pixel-level predictions. Together, these components allow the model to more accurately distinguish between ambiguous regions and reduce misclassification errors caused by overlapping semantics. As shown in Tab. 1, our method achieves superior mIoU performance compared to a range of recent state-of-the-art methods, including both convolution-based and transformer-based backbones. Notably, CPC outperforms strong baselines such as ToCo [59] and SFC [62], demonstrating the effectiveness of our dual-level learning strategy in producing high-quality pseudo labels.
@@
Model Pub. Backbone Val USAGE [60] ICCV23 ResNet38 71.9 SAS [63] AAAI23 ViT-B/16 69.5 MCTformer [33] CVPR22 DeiT-S 61.7 SIPE [26] CVPR22 ResNet50 58.6 ViT-PCM [10] ECCV22 ViT-B/16 69.3 AFA [34] CVPR22 MiT-B1 63.8 ToCo [59] CVPR23 ViT-B/16 70.5 TSCD [64] AAAI23 MiT-B1 67.3 FPR [61] ICCV23 ResNet38 70.0 SFC [62] AAAI24 ViT-B/16 71.2 IACD [2] ICASSP24 ViT-B/16 71.4 APC [1] EAAI25 ViT-B/16 72.3 CPC Ours ViT-B/16 73.4
@@
Model Pub. Backbone mIoU (%) MCTformer [33] CVPR22 Resnet38 42.0 ViT-PCM [10] ECCV22 ViT-B/16 45.0 SIPE [26] CVPR22 Resnet38 43.6 TSCD [64] AAAI23 MiT-B1 40.1 SAS [63] AAAI23 ViT-B/16 44.5 FPR [61] ICCV23 ResNet38 43.9 ToCo [59] CVPR23 ViT-B 42.3 SFC [62] AAAI24 ViT-B/16 44.6 IACD [2] ICASSP24 ViT-B/16 44.3 PGSD [65] TCSVT24 ViT-B/16 43.9 APC [1] EAAI25 ViT-B/16 45.7 CPC Ours ViT-B/16 46.7
Semantic Segmentation Results. To evaluate the effectiveness of our proposed pseudo-label generation strategy, we adopt DeepLab V2 [52] as the segmentation backbone and train it using the pseudo labels generated by our CPC framework. We report semantic segmentation performance on the PASCAL VOC 2012 validation set and compare it with recent state-of-the-art WSSS approaches, as presented in Tab. 2. Our CPC framework yields superior segmentation performance, achieving an mIoU of 72.2% and outperforming prior transformer-based methods such as ToCo [59], SFC [62], and IACD [2]. These results confirm the advantage of combining inter-class category clustering with intra-class contrastive learning for generating high-quality pseudo labels. The enhanced quality of pseudo labels leads to clearer delineation of complex object structures and reduced confusion between visually similar classes.